If length of a rectangle is halved and breadth is doubled then the area of the rectangle obtained remains same.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement isTrue.
We know that, area of rectangle = length × breadth
As per the condition given in the question, length of a rectangle is halved = 1⁄2

= l/2

Breadth is doubled = 2b Then, new area = l/2 × 2b

= l × b.
Therefore, length of a rectangle is halved and breadth is doubled then the area of the rectangle obtained remains same.


Thanks